# Descent spectral sequences through synthetic spectra

### v1

@@ Introduction
Theorem C (Cors. 4.13 and 4.15).
The signatures of the synthetic $\mathbf{E}_{\infty}$-rings $\mathrm{SMF}\vcentcolon=\mathcal{O}^{\mathrm{syn}}(\mathsf{M}_{\mathrm{Ell}})$ and $\mathrm{Smf}\vcentcolon=\mathcal{O}^{\mathrm{syn}}(\overline{\mathsf{M}}_{\mathrm{Ell}})$ are the DSS for $\mathrm{TMF}$ and $\mathrm{Tmf}$, respectively. There is an equivalence of synthetic $\mathbf{E}_{\infty}$-rings $\mathrm{SMF}\cong\operatorname{\nu}\mathrm{TMF}$, but the synthetic spectra $\mathrm{Smf}$ and $\operatorname{\nu}\mathrm{Tmf}$ are not equivalent.
The equivalence $\mathrm{SMF}\cong\operatorname{\nu}\mathrm{TMF}$ follows from Th. A, and recovers the well-known result that the ANSS for $\mathrm{TMF}$ is isomorphic to its DSS. For $\mathrm{Tmf}$, these spectral sequences differ. In fact, we can quantify the difference between $\mathrm{Smf}$ and $\operatorname{\nu}\mathrm{Tmf}$; see Cor. 4.15 for a detailed comparison.
The framework of Th. B allows the fruits of synthetic spectra to be brought to bear upon DSSs. For example, one can now think of the DSS for $\mathrm{Tmf}$ as a modified Adams–Novikov spectral sequence for $\mathrm{Tmf}$. As a result, tools such as localisations, power operations, and Toda brackets, along with versions of Moss’s theorem, may be used more naturally in this context. These ideas will be used in [CDN24] to calculate the homotopy groups of $\mathrm{Tmf}$ whilst avoiding the circular logic in the current literature.
@@
In § 4, we apply the above theory in the two examples mentioned in the introduction. First, in § 4.1, we consider the fixed points of the action of a finite group $G$ on the a Landweber exact $\mathbf{E}_{\infty}$-ring. This includes the action of a finite subgroup of the Morava stabiliser group on Morava E-theory, and also the case of K-theory mentioned in the introduction. Second, we study topological modular forms in § 4.2. We introduce the synthetic spectra $\mathrm{SMF}$ and $\mathrm{Smf}$ in Df. 4.8, and then show that $\operatorname{\nu}\mathrm{TMF}\cong\mathrm{SMF}$ and $\operatorname{\nu}\mathrm{Tmf}\ncong\mathrm{Smf}$, proving Th. C.
@@ Acknowledgements
We thank Shaul Barkan, Sil Linskens, Lennart Meier, and Lucas Piessevaux for their stimulating discussions around these topics, and Lennart Meier for his feedback on earlier drafts of this article.
@@ 1   Synthetic spectra and spectral sequences
For this section, let $E$ be a homotopy associative ring spectrum of Adams type ([Pst23, Df.3.14]), and write $\mathrm{Syn}_{E}$ for the $\infty$-category of $E$-based synthetic spectra of [Pst23]. For an introduction to synthetic spectra, we refer to [Pst23] and [BHS23, §9], and we further follow much of the notation from [CD24, §1–2].
@@ 1.1   The synthetic analogue functor
Let $M$ be a spectrum that admits the structure of a homotopy $E$-module, meaning an $E$-module object in $\mathrm{h}\mathrm{Sp}$.
@@
    We have a natural isomorphism of bigraded abelian groups
@@
    For any spectrum $X$, the natural map
    $\nu M\otimes\nu X\to\nu(M\otimes X)$
    coming from the lax monoidal structure on $\nu$ is an equivalence.
@@
The first item is [Pst23, Pr.4.60]. For the second item, note that this map is a $\nu E$-equivalence by [Pst23, Lm.4.24]. However, both the source and target are $\nu E$-local. Indeed, using the lax monoidal structure on $\nu$ from [Pst23, Lm.4.4], we see that they both admit the structure of a homotopy $\nu E$-module, and in particular are $\nu E$-local. ∎
@@
As in most of this paper, consider the case where $E=\mathrm{MU}$, so that $\mathrm{Syn}_{\mathrm{MU}}$ is the categorification of the Adams–Novikov spectral sequence. If $R$ is a complex-orientable ring spectrum, then $R$ in particular admits the structure of a homotopy $\mathrm{MU}$-module. Hence the above results apply to $\nu R$. In particular, this applies to all Landweber-exact ring spectra.
@@ 1.2   Filtered spectra
$\sigma(X)^{n}=\operatorname{map}_{\mathrm{Syn}_{E}}(\mathbf{S},\,\Sigma^{0,-n}X)\xrightarrow{\operatorname{map}(\mathbf{1},\tau)}\operatorname{map}_{\mathrm{Syn}_{E}}(\mathbf{S},\,\Sigma^{0,-n+1}X)=\sigma(X)^{n-1}.$
The functor $\sigma$ comes equipped with a natural lax symmetric monoidal structure: see [CD24, Th.1.4(4)], where the notation $\Gamma$ is used for $\sigma$.
@@
Let $X$ be a synthetic spectrum. Then the signature of $X$ is the spectral sequence associated to the filtered spectrum $\sigma X$.
The functor $\sigma$ is a right adjoint [^4] The left adjoint to $\sigma$ is the universal extension of the filtered object $\mathbf{Z}\to\mathrm{Syn}_{E}$ defined by the $\tau$-tower of the unit. and, in many cases, part of a monadic adjunction [Law24, Cor.6.1]. For now, we will just need the following.
@@ 1.3   Bousfield–Kan spectral sequences
In particular, if $X^{\bullet}$ is a cosimplicial spectrum, then the BKSS computing the homotopy groups of the limit of $X^{\bullet}$ is the spectral sequence associated with the filtered spectrum $\mathrm{D\acute{e}c}(X^{\bullet})$. If the resolution $X^{\bullet}$ is nice enough, then we can package this BKSS as a synthetic spectrum, or more precisely as a synthetic lift of $\lim X^{\bullet}$. Moreover, this synthetic spectrum recovers the filtered spectrum $\mathrm{D\acute{e}c}(X^{\bullet})$, together with its multiplicative structure.
@@ 1.4   The signature of a synthetic analogue
For most of this article, we find ourselves in the situation where $E$ has an $\mathbf{E}_{\infty}$-ring structure: our comparison results happen in the case where $E=\mathrm{MU}$. For the sake of completeness, we will in this section compare this classical definition of the Adams filtration with the modern definition $\sigma\circ\nu$, and show that the monoidal structures also agree. This is certainly well known: see [Pst23, Rmks.4.62–65] and [BHS23, §A.1], but we include it for completeness and to establish notation to be used later.
@@ 2   The synthetic lift of an even-periodic refinement
If $\mathsf{X}$ is a stack equipped with a sheaf of spectra $\mathcal{F}$, then the descent spectral sequence (DSS) for $\mathcal{F}$ attempts to compute the homotopy groups of the global sections $\Gamma(X,\mathcal{F})$ by means of the homotopy sheaves $\pi_{*}\mathcal{F}$. We study the DSS from a synthetic point of view in the context of an even-periodic refinement of a classical Deligne–Mumford stack. We show that, in this case, the DSS is most naturally viewed as the signature of a certain $\mathrm{MU}$-synthetic spectrum. First, we define a filtered model for the DSS in § 2.1, and then construct this $\mathrm{MU}$-synthetic spectrum in § 2.2.
@@ 2.2   Descent spectral sequences in synthetic spectra
Let $E$ be a complex-oriented $\mathbf{E}_{\infty}$-ring, with an action of a finite group $G$. Then the quotient $\operatorname{Spec}(E)/G$ represents an even-periodic refinement of $\operatorname{Spec}(\pi_{0}E)/G$. Common examples include the Morava E-theories $\mathrm{E}_{n}$ and the action by various finite subgroups of the Morava stabiliser group. An integral example of this form is $\operatorname{Spec}(\mathrm{KU})/C_{2}$ from the introduction.
@@
Let $\overline{\mathsf{M}}_{\mathrm{Ell}}$ be the compactification of the moduli stack of elliptic curves. Then the sheaf $\mathcal{O}^{\mathrm{top}}$ is an even-periodic refinement of $\overline{\mathsf{M}}_{\mathrm{Ell}}$ by design; see [Goe10]. More general examples of this flavour include anything coming out of Lurie’s theorem, see [Dav20, §5], such as the spectra of topological automorphic forms of Behrens–Lawson [BL10].
@@
$\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(A)=\nu\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(A)\to\lim\left(\nu\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(B^{0})\Rightarrow\nu\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(B^{1})\Rrightarrow\cdots\right)$ (2.9)
@@
Being $\tau$-complete, these synthetic spectra are to some extent fully captured by their signature (or more precisely, their associated filtered spectra). We can compute their signature using the results in § 1; this gives us Th. B.
@@
In particular, the signature of $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$ is naturally equivalent to the DSS for $\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(\mathsf{X})$.
@@
Remark 2.13.
The equivalence above is also natural in $\mathsf{X}$. More specifically, if $f\colon\mathsf{Y}\to\mathsf{X}$ is a morphism of Deligne–Mumford stacks, then there is a commutative diagram of filtered spectra
[…]
induced from the base-change along $f$ of an affine étale cover $\operatorname{Spec}A\to\mathsf{X}$. Indeed, by Th. 1.18, we have a natural monoidal equivalence
$\mathrm{DSS}(\mathsf{X},\mathcal{O}^{\mathrm{top}}_{\mathsf{X}})=\mathrm{D\acute{e}c}(\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(\operatorname{Spec}A^{\bullet}))\cong\sigma\lim(\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\operatorname{Spec}A^{\bullet}))=\operatorname{\sigma}\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$
where $\operatorname{Spec}A\to\mathsf{X}$ is an étale cover.
Further, Pr. 1.6 tells us that $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}$ is levelwise a synthetic lift of $\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}$. We can be more precise about their relationship, as follows. In the following proposition, we use the notation $\nu\circ\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}$ to denote the composition of functors; this is not to be confused with the sheafification thereof, which we denote by $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}$.
@@ 3.2   Identifying the abutment
These affineness results do not hold synthetically, even in the simplest examples. Consider, as in the introduction, the even-periodic refinement $\operatorname{Spec}(\mathrm{KU})/C_{2}$ of the affine map $BC_{2}\to\mathsf{M}_{\mathrm{FG}}$. The corresponding $\infty$-category of quasicoherent synthetic sheaves may be identified with $\mathrm{Mod}_{\operatorname{\nu}\mathrm{KU}}(\mathrm{Syn}_{\mathrm{MU}})^{hC_{2}}$, as in [MM15, Pr.2.16]. We claim that the global sections functor
@@
$\mathcal{G}(\mathcal{O}^{\mathrm{top}}_{\mathsf{X}})\cong f^{\ast}q_{\ast}\mathcal{O}_{\operatorname{Spec}L}\cong\bar{q}_{*}\bar{f}^{*}\mathcal{O}_{\operatorname{Spec}L}\cong\bar{q}_{\ast}\bar{q}^{\ast}\omega^{\otimes\ast/2}.$

### v2

@@ Introduction
Theorem C (Cors. 4.15 and 4.17).
The signature spectral sequences of the synthetic $\mathbf{E}_{\infty}$-rings $\mathrm{SMF}\vcentcolon=\mathcal{O}^{\mathrm{syn}}(\mathsf{M}_{\mathrm{Ell}})$ and $\mathrm{Smf}\vcentcolon=\mathcal{O}^{\mathrm{syn}}(\overline{\mathsf{M}}_{\mathrm{Ell}})$ are the DSS for $\mathrm{TMF}$ and $\mathrm{Tmf}$, respectively. There is an equivalence of synthetic $\mathbf{E}_{\infty}$-rings $\mathrm{SMF}\cong\operatorname{\nu}\mathrm{TMF}$, but the synthetic spectra $\mathrm{Smf}$ and $\operatorname{\nu}\mathrm{Tmf}$ are not equivalent.
The equivalence $\mathrm{SMF}\cong\operatorname{\nu}\mathrm{TMF}$ follows from Th. A, and recovers the well-known result that the ANSS for $\mathrm{TMF}$ is isomorphic to its DSS. For $\mathrm{Tmf}$, these spectral sequences differ. In fact, we can quantify the difference between $\mathrm{Smf}$ and $\operatorname{\nu}\mathrm{Tmf}$; see Cor. 4.17 for a detailed comparison.
The framework of Th. B allows the fruits of synthetic spectra to be brought to bear upon DSSs. For example, one can now think of the DSS for $\mathrm{Tmf}$ as a modified Adams–Novikov spectral sequence for $\mathrm{Tmf}$. As a result, tools such as localisations, power operations, and Toda brackets, along with versions of Moss’s theorem, may be used more naturally in this context. These ideas are used in [CDvN24] to calculate the homotopy groups of $\mathrm{Tmf}$ whilst avoiding the circular logic in the current literature. In particular, we highlight that the ANSS for $\mathrm{Tmf}$ is unmanagable and otherwise work with the synthetic spectrum $\mathrm{Smf}$; see [CDvN24, §1.2.1].
@@
In § 4, we apply the above theory in the two examples mentioned in the introduction. First, in § 4.1, we consider the fixed-points of the action of a finite group $G$ on the a Landweber exact $\mathbf{E}_{\infty}$-ring. This includes the action of a finite subgroup of the Morava stabiliser group on Morava E-theory, and also the case of K-theory mentioned in the introduction. Second, we study topological modular forms in § 4.2. We introduce the synthetic spectra $\mathrm{SMF}$ and $\mathrm{Smf}$ in Df. 4.10, and then show that $\operatorname{\nu}\mathrm{TMF}\cong\mathrm{SMF}$ and $\operatorname{\nu}\mathrm{Tmf}\ncong\mathrm{Smf}$, proving Th. C.
@@ Acknowledgements
We thank Shaul Barkan, Sil Linskens, Lennart Meier, and Lucas Piessevaux for their stimulating discussions around these topics, and Lennart Meier for his feedback on earlier drafts of this article. Thank you as well to the anonymous referee for their suggestions.
@@ 1   Synthetic spectra and spectral sequences
For this section, let $E$ be a homotopy associative ring spectrum of Adams type ([Pst23, Df.3.14]), and write $\mathrm{Syn}_{E}$ for the $\infty$-category of $E$-based synthetic spectra of [Pst23]. For an introduction to synthetic spectra, we refer to [Pst23] and [BHS23, §9], and we further follow much of the notation from [CD24, §2–3].
@@ 1.1   The synthetic analogue functor
    Let $M$ be a spectrum that admits the structure of a homotopy $E$-module. Then we have a natural isomorphism of bigraded abelian groups
@@
    Let $X$ be spectrum that can be written as a filtered colimit of finite $E$-projective spectra. Then for any spectrum $Y$, the natural map
    $\nu X\otimes\nu Y\to\nu(X\otimes Y)$
    coming from the lax monoidal structure on $\nu$ is an equivalence. In particular, this holds if $X=E$.
@@
The first item is [Pst23, Pr.4.60], and the second is [Pst23, Lm.4.24]. ∎
@@
As in most of this paper, consider the case where $E=\mathrm{MU}$, so that $\mathrm{Syn}_{\mathrm{MU}}$ is the categorification of the Adams–Novikov spectral sequence. If $R$ is a Landweber-exact ring spectrum, then $R$ can be written as a filtered colimit of finite even spectra; see [HS99, §2.1]. A finite even spectrum in particular has projective $\mathrm{MU}_{*}$-homology, so Pr. 1.3(2) applies.
@@ 1.2   Filtered spectra
$\sigma(X)^{n}=\operatorname{map}_{\mathrm{Syn}_{E}}(\mathbf{S},\,\Sigma^{0,-n}X)\xrightarrow{\operatorname{map}(\mathbf{S},\tau)}\operatorname{map}_{\mathrm{Syn}_{E}}(\mathbf{S},\,\Sigma^{0,-n+1}X)=\sigma(X)^{n-1}.$
The functor $\sigma$ comes equipped with a natural lax symmetric monoidal structure: see [CD24, Th.2.5(4)], where the notation $\Gamma$ is used for $\sigma$.
@@
Let $X$ be a synthetic spectrum. Then the signature of $X$ is the filtered spectrum $\sigma X$ and the associated spectral sequence is called the signature spectral sequence.
The functor $\sigma$ is a right adjoint [^4] The left adjoint to $\sigma$ is the universal extension of the functor $\mathbf{Z}\to\mathrm{Syn}_{E}$ defined by the $\tau$-tower of the unit. and, in many cases, part of a monadic adjunction [Law24, Cor.6.1]. For now, we will just need the following.
@@ 1.3   Bousfield–Kan spectral sequences
In particular, if $X^{\bullet}$ is a cosimplicial spectrum, then the BKSS computing the homotopy groups of the limit of $X^{\bullet}$ is the spectral sequence associated with the filtered spectrum $\mathrm{D\acute{e}c}(X^{\bullet})$. If the resolution $X^{\bullet}$ is nice enough, then we can package this BKSS as a synthetic spectrum, or more precisely as a synthetic lift of $\lim X^{\bullet}$. Moreover, this synthetic spectrum recovers the filtered spectrum $\mathrm{D\acute{e}c}(X^{\bullet})$, together with its multiplicative structure. A similar results appears in [BP21, Lm.9.18].
@@ 1.4   The signature of a synthetic analogue
For most of this article, we find ourselves in the situation where $E$ has an $\mathbf{E}_{\infty}$-ring structure: our comparison results happen in the case where $E=\mathrm{MU}$. For the sake of completeness, we will in this section compare this classical definition of the Adams filtration with the modern definition $\sigma\circ\nu$, and show that the monoidal structures also agree. This is certainly well known: see [Pst23, Rmks.4.62–65], [BHS23, §A.1], and [PP21, Df.5.58, Th.5.60], but we include it for completeness and to establish notation to be used later.
@@ 2   The synthetic lift of an even-periodic refinement
If $\mathsf{X}$ is a stack equipped with a sheaf of spectra $\mathcal{F}$, then the descent spectral sequence (DSS) for $\mathcal{F}$ attempts to compute the homotopy groups of the global sections $\Gamma(X,\mathcal{F})$ by means of the homotopy sheaves $\pi_{*}\mathcal{F}$. We study the DSS from a synthetic point of view in the context of an even-periodic refinement of a classical Deligne–Mumford stack. We show that, in this case, the DSS is most naturally viewed as the signature spectral sequence of a certain $\mathrm{MU}$-synthetic spectrum. First, we define a filtered model for the DSS in § 2.1, and then construct this $\mathrm{MU}$-synthetic spectrum in § 2.2.
@@ 2.2   Descent spectral sequences in synthetic spectra
Let $E$ be an even-periodic complex-oriented $\mathbf{E}_{\infty}$-ring, with an action of a finite group $G$. Then the quotient $\operatorname{Spec}(E)/G$ represents an even-periodic refinement of $\operatorname{Spec}(\pi_{0}E)/G$. Common examples include the Morava E-theories $\mathrm{E}_{n}$ and the action by various finite subgroups of the Morava stabiliser group. An integral example of this form is $\operatorname{Spec}(\mathrm{KU})/C_{2}$ from the introduction.
@@
Let $\overline{\mathsf{M}}_{\mathrm{Ell}}$ be the compactification of the moduli stack of elliptic curves. Then the Goerss–Hopkins–Miller sheaf $\mathcal{O}^{\mathrm{top}}$ is an even-periodic refinement of $\overline{\mathsf{M}}_{\mathrm{Ell}}$ by design; see [Goe10] and [DL25, Th.B]. More general examples of this flavour include anything coming out of Lurie’s theorem, see [Dav24, §5], such as the spectra of topological automorphic forms of Behrens–Lawson [BL10].
@@
$\nu\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(A)\to\lim\left(\nu\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(B^{0})\Rightarrow\nu\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(B^{1})\Rrightarrow\cdots\right)$ (2.9)
@@
Being $\tau$-complete, these synthetic spectra are to some extent fully captured by their signature. We can compute their signature using the results in § 1; this gives us Th. B.
@@
In particular, the signature spectral sequence of $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$ is naturally equivalent to the DSS for $\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(\mathsf{X})$.
A more general version of this theorem for quasi-coherent sheaves over an even-periodic refinement appears in [Nig25].
@@
The following is a computationally useful corollary of the above fact. It states that the signature of $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$ always satisfies a checkerboard pattern on its $\mathrm{E}_{2}$-page.
Corollary 2.13.
For an even-periodic refinement $(\mathsf{X},\mathcal{O}^{\mathrm{top}}_{\mathsf{X}})$, the bigraded homotopy groups of $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$ are even, in the sense that their bigraded homotopy groups vanish when the sum of the two bidegrees is odd.
In [Nig25], this is related to even $\mathrm{MU}$-synthetic spectra as in [Pst23, Df.5.8].
Proof.
By Cor. 2.12, the mod $\tau$ homotopy groups of $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$ give the $\mathrm{E}_{2}$-page of the DSS for $\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(\mathsf{X})$. As $\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}$ is an even-periodic refinement, this spectral sequence is concentrated in such degrees. ∎
Furthermore, Pr. 1.6 tells us that $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}$ is levelwise a synthetic lift of $\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}$. We can be more precise about their relationship, as follows. In the following proposition, we use the notation $\nu\circ\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}$ to denote the composition of functors; this is not to be confused with the sheafification thereof, which we denote by $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}$.
@@ 3.2   Identifying the abutment
These affineness results fail synthetically, even in the simplest examples. Consider, as in the introduction, the even-periodic refinement $\operatorname{Spec}(\mathrm{KU})/C_{2}$ of the affine map $BC_{2}\to\mathsf{M}_{\mathrm{FG}}$. The corresponding $\infty$-category of quasicoherent synthetic sheaves may be identified with $\mathrm{Mod}_{\operatorname{\nu}\mathrm{KU}}(\mathrm{Syn}_{\mathrm{MU}})^{hC_{2}}$, as in [MM15, Pr.2.16]. We claim that the global sections functor
@@
The key problem occurring in the above remarks is one of hypercompleteness: if we consider the synthetic global sections functor taking values in the $\infty$-category of modules over $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$ in hypercomplete $\mathrm{MU}$-synthetic spectra, then one does obtain various cocontinuity statements; see [Nig25] for more.
@@
$\mathcal{G}(\mathcal{O}^{\mathrm{top}}_{\mathsf{X}})\cong f^{\ast}q_{\ast}\mathcal{O}_{\operatorname{Spec}L}\cong\bar{q}_{*}\bar{f}^{*}\mathcal{O}_{\operatorname{Spec}L}\cong\bar{q}_{\ast}\bar{q}^{\ast}\omega_{\mathsf{X}}^{\otimes\ast/2}.$